OpenStudy (anonymous):

In this problem, what are the steps I need to use? Please help. (8x+5)(2x^2+9x+5)

OpenStudy (anonymous):

(8x+5)(2x^2+9x+5)= 18x^3+72x^2+40x+10x^2+45x+25= 18x^3+82x^2+85x+25

OpenStudy (anonymous):

are these the steps I need to take? or is it the answer?

myininaya (myininaya):

\[(8x+5)(2x^2+9x+5)=8x(2x^2+9x+5)+5(2x^2+9x+5)\]

myininaya (myininaya):

\[=8x(2x^2)+8x(9x)+8x(5)+5(2x^2)+5(9x)+5(5)\]

myininaya (myininaya):

Do you see what I did so far?

OpenStudy (anonymous):

I do.

myininaya (myininaya):

Great! Now it is just matter of using law of exponents and the fact that multiplication is commutative

OpenStudy (anonymous):

Yes. It is the answer

myininaya (myininaya):

and also associative \[=8(2)x(x^2)+8(9)x(x)+8(5)x+6(2)x^2+5(9)(x)+5(5)\]

OpenStudy (anonymous):

oh ok...

myininaya (myininaya):

\[=16x^{2+1}+72x^{1+1}+40x+12x^2+45x+25\]

OpenStudy (anonymous):

Oh wow, that last one looks confusing.

myininaya (myininaya):

You don't have to do that step I was just trying to use some basic rules

myininaya (myininaya):

\[=16x^3+72x^2+40x+12x^2+45x+25\] Combine like terms and you will be done

myininaya (myininaya):

But if you need me to go over that step with you again I will

myininaya (myininaya):

\[=8x(2x^2)+8x(9x)+8x(5)+5(2x^2)+5(9x)+5(5) \] This was the expression I had before I made it confusing for you I think right?

OpenStudy (anonymous):

so we are combining like terms for the last step?

myininaya (myininaya):

Yes! :)

OpenStudy (anonymous):

Yes that was before it got confusing.

OpenStudy (anonymous):

Ok so let me give this a shot then.

myininaya (myininaya):

Ok...I will give you some time to yourself and say something when you want me to come back k?

OpenStudy (anonymous):

which one are we combining like terms for so I can make sure I have the right one wrote down.

myininaya (myininaya):

\[=16x^3+72x^2+40x+12x^2+45x+25 \]

OpenStudy (anonymous):

Ok i had the wrong one...give me a minute let me see what I get..

OpenStudy (anonymous):

I am not so sure I got it right, but here is what I got. 185x^9+25

myininaya (myininaya):

Oh no that is not right

OpenStudy (anonymous):

Ok I didn't think so. I am not good at this stuff...

myininaya (myininaya):

like terms are terms with the same variable part

OpenStudy (anonymous):

Ok by variable you mean the "x" right?

myininaya (myininaya):

You want to look for all the terms with x^3 (and add those together) Then you want to look for all terms with x^2 (and add those together) Then you want to look for all terms with x (and add those together) And add any constant terms that are remaining together

OpenStudy (anonymous):

Oh oK...trying again.

myininaya (myininaya):

Ok go you! You can do it! :)

OpenStudy (anonymous):

Ok trying again.

OpenStudy (anonymous):

16x^3+84x^2+85x+25

myininaya (myininaya):

+10000 points :) Good job!

OpenStudy (anonymous):

THANK YOU for showing me how to work it out!!!!!

myininaya (myininaya):

But wait

myininaya (myininaya):

I think I made a typing error above k?

OpenStudy (anonymous):

ok...

myininaya (myininaya):

\[=8x(2x^2)+8x(9x)+8x(5)+5(2x^2)+5(9x)+5(5) \] We had this and I accidentally wrote \[=8(2)x(x^2)+8(9)x(x)+8(5)x+6(2)x^2+5(9)(x)+5(5) \] See how that one number changed from a 5 to a 6 my bad :(

myininaya (myininaya):

\[=16x^3+72x^2+40x+10x^2+45x+25\]

OpenStudy (anonymous):

Ok so need to try it again right? will make the answer different.

myininaya (myininaya):

yes it will and sorry about that

OpenStudy (anonymous):

Not a problem it happens. I got 16x^3+82x^2+85x+25

myininaya (myininaya):

Bingo! Now you can take the points and be happy. lol

OpenStudy (anonymous):

Thank You so much! Your a lot more patient than some of these people on here, and I like being showed how its worked out and not having the answer just given to me! Thanks again!!

myininaya (myininaya):

You are totally welcome! :) I like people who want to learn the material so you are really awesome! Go you! Have a good night!

OpenStudy (anonymous):

You too!